# Quantum Algorithms in a Superposition of Spacetimes

### v1

@@ 2.1 Computable Order Interference and Quantum State Generation
When we think of the ability to generate new quantum states, there comes to mind one of the most fundamental hard problems in computer science – the Statistical Difference ($\mathsf{SD}$) problem [SV03]. The input to the $\mathsf{SD}$ problem is a pair of classical circuits $\left(C_{0},C_{1}\right)$, and we consider their output distributions (i.e., the output distribution for a uniformly random input string). The problem is to decide the total variation distance (also named statistical distance) between these two distributions. Formally, for functions $a,b:\mathbb{N}\rightarrow[0,1]$ such that $\forall n\in\mathbb{N}:a(n)\leq b(n)$, an input to the $\mathsf{SD}_{a(n),b(n)}$ problem is a pair $\left(C_{0},C_{1}\right)$ of classical circuits with the same output size, $C_{b}:\{0,1\}^{k_{b}}\rightarrow\{0,1\}^{k}$ for some $k_{0},k_{1},k\in\mathbb{N}$. We denote the problem input size with $n:=|C_{0}|+|C_{1}|$ and think of the circuits as a pair of sampler circuits, that given uniformly random $k_{b}$ bits sample a $k$-bit output. The problem comes with the promise that it is either the case that the statistical distance between the output distributions of the circuits is at most $a(n)$, in which case we denote $\left(C_{0},C_{1}\right)\in\prod_{\text{YES}}$, or the statistical distance is more than $b(n)$, in which case we denote $\left(C_{0},C_{1}\right)\in\prod_{\text{NO}}$. Generally speaking, the larger the gap $b(n)-a(n)$, the easier the problem. In this work, our focus is on the family of problems $\mathsf{SD}_{a(n),b(n)}$, for functions $a(n),b(n)$ such that there exists a polynomial $\mathsf{poly}(n)$ with $b(n)-2a(n)+a(n)^{2}\geq\frac{1}{\mathsf{poly}(n)}$  [^5] The choice of $b(n)^{2}-2a(n)+a(n)^{2}$ comes from the relation between total variation distance and fidelity, that is, if fidelity between the output distributions of the circuits was the measure determining $a(n)$, $b(n)$, then we will simply be interested in $b(n)-a(n)\geq\frac{1}{\mathsf{poly}(n)}$.. From hereon we drop the subscript $a(n)$, $b(n)$ from $\mathsf{SD}_{a(n),b(n)}$ and write $\mathsf{SD}_{\mathsf{poly}}$ to implicitly refer to the family of problems $\mathsf{SD}_{a(n),b(n)}$ for such functions $a(n)$, $b(n)$. The $\mathsf{SD}_{\mathsf{poly}}$ problem is a notoriously hard problem in computer science (also for quantum computers), and it is at least as hard as some of the widely studied computational problems in cryptography and in combinatorics, some examples include: Quadratic Residuosity, Lattice Isomorphism, Discrete Logarithm, Decisional Diffie-Hellman, Graph Isomorphism, the Gap Closest Vector Problem and the Gap Shortest Vector Problem.
The thing that connects $\mathsf{SD}_{\mathsf{poly}}$ and the ability to generate non-trivial quantum states is a quantum procedure called the Swap Test [${}^{+}$97, BCWDW01]. It is observed in [ATS03] that the $\mathsf{SD}_{\mathsf{poly}}$ problem could be efficiently solved if we had the ability to efficiently generate the output distribution states of classical circuits. If given input $\left(C_{0},C_{1}\right)$ we could efficiently generate the normalized $k$-qubit quantum states,
@@
then we can execute a Swap Test on the pair of states $|{C_{0}\left(R\right)}\rangle$, $|{C_{1}\left(R\right)}\rangle$, which approximates (with some inverse polynomial error) the inner product between the states as vectors. Since the inner product between these states relates to the statistical distance between the output distributions of the circuits $C_{0}$, $C_{1}$, a Swap Test lets us estimate the statistical distance. Finally, as long as the gap $b(n)^{2}-2a(n)+a(n)^{2}$ is at least inverse polynomial, the error in the approximation of the Swap Test can be overcome and we can decide whether the pair of circuits is in $\prod_{\text{YES}}$ or $\prod_{\text{NO}}$. So, the $\mathsf{SD}_{\mathsf{poly}}$ problem can be reduced to the quantum state generation of $|{C_{0}\left(R\right)}\rangle$, $|{C_{1}\left(R\right)}\rangle$. With accordance to the hardness of the $\mathsf{SD}_{\mathsf{poly}}$ problem, there is no known way for a quantum computer to generate in polynomial time the output distribution state for a general classical circuit $C$.
@@ 2.2 The Sequentially Invertible Statistical Difference Problem
It is a standard fact in quantum computing that for an efficient classical circuit $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ there is an $\left(k^{\prime}+k\right)$-qubit unitary $U_{x,C}$, of roughly the same computational complexity as $C$, that maps: $\forall x\in\{0,1\}^{k^{\prime}},y\in\{0,1\}^{k}:U_{x,C}\cdot|{x,y}\rangle=|{x,y\oplus C(x)}\rangle$. It is in particular easy to quantumly generate the $\left(k^{\prime}+k\right)$-qubit state
@@
Sequentially invertible functions.
@@
Our next idea is to parameterize invertibility. Our parameterization technique is not necessarily related to quantum computation and may be of independent interest. Note that that sampler circuits are equivalent to randomized circuits, which take the form $C:\{0,1\}^{k}\times\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$, where the randomness $z\in\{0,1\}^{k^{\prime}}$ is only considered as auxiliary input, and together with it there’s a main input $x\in\{0,1\}^{k}$. We can thus consider sampler and randomized circuits interchangeably.
@@
Our definition of parameterized invertibility of a function is as follows. Let $f:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ a function. For $r\leq k^{\prime}$, we say that $f$ is $r$-sequentially invertible if there is an $r$-invertible function sequence that samples from the output distribution of $f$ [^6] The full definition of $\left(r,t,\ell,\varepsilon\right)$-sequential invertibility, which takes into consideration more parameters and not only $r$, is given in Section 4.3. For the needs for this overview, we give the minimal definitions.. An $r$-invertible function sequence is a sequence of pairs of functions $\left(f_{i,\rightarrow},f_{i,\leftarrow}\right)_{i\in[\ell]}$ such that,
  • There exists $\bar{k}\in\mathbb{N}$ such that all functions have input and output size $\bar{k}$, and randomness bounded by $r$. Additionally, each pair of opposite-direction functions use the same amount of randomness. Formally, for every $i\in[\ell]$ there exists $r_{i}\leq r$ such that
    $\forall Y\in\{\rightarrow,\leftarrow\}:f_{i,Y}:\{0,1\}^{\bar{k}}\times\{0,1\}^{r_{i}}\rightarrow\{0,1\}^{\bar{k}}\enspace.$
  • The function directions are inverses of each other, per hard-wired randomness, that is:
    $\forall i\in[\ell],z\in\{0,1\}^{r_{i}},x\in\{0,1\}^{\bar{k}}:x=f_{i,\leftarrow}\left(f_{i,\rightarrow}\left(x;z\right);z\right)\enspace.$
@@
We say that the sequence samples from $f$, if $\bar{k}=k$, and we have the following equivalence of distributions,
$\Bigl{\{}f(z)\;|\;z\leftarrow\{0,1\}^{k^{\prime}}\Bigr{\}}\equiv$
@@
$\biggl{\{}f_{\ell,\rightarrow}\left(\cdots f_{2,\rightarrow}\left(f_{1,\rightarrow}\left(0^{\bar{k}};z_{1}\right);z_{2}\right)\cdots;z_{\ell}\right)\;\Big{|}\;z_{1}\leftarrow\{0,1\}^{r_{1}},\cdots,z_{\ell}\leftarrow\{0,1\}^{r_{\ell}}\biggr{\}}\enspace.$
@@
The intuition is that the smaller the parameter $r$ is, ”the more the function is invertible”. We give a few examples. For any invertible function $f:\{0,1\}^{k}\rightarrow\{0,1\}^{k}$, $f$ is 00-sequentially invertible (for $\ell=1$, that is, by a single pair $\left(f_{1,\rightarrow},f_{1,\leftarrow}\right)$). $f_{\oplus}:\{0,1\}^{2\cdot k}\rightarrow\{0,1\}^{k}$, that for two strings $a,b\in\{0,1\}^{k}$ outputs their bit-wise logical XOR: $f_{\oplus}(a,b):=a\oplus b$, is $1$-sequentially invertible (by a sequence of $\ell=k$ pairs), and $f_{\land}:\{0,1\}^{2\cdot k}\rightarrow\{0,1\}^{k}$, that for two strings $a,b\in\{0,1\}^{k}$ outputs their bit-wise logical AND: $f_{\land}(a,b):=a\land b$, is $2$-sequentially invertible (again, by a sequence of $k$ pairs). We leave it as an instructive exercise for the reader to verify these claims.
@@
where for each $b\in\{0,1\}$, $\left(C^{b}_{i,\rightarrow},C^{b}_{i,\leftarrow}\right)_{i\in[\ell]}$ is an $r(n)$-invertible circuit sequence, where $n$ is the input size,
$n:=\bigg{|}\left(\left(C^{0}_{i,\rightarrow},C^{0}_{i,\leftarrow}\right)_{i\in[\ell]},\left(C^{1}_{i,\rightarrow},C^{1}_{i,\leftarrow}\right)_{i\in[\ell]}\right)\bigg{|}\enspace,$
and such that both sequences have the same input and output size $k$, that is, for every $i\in[\ell]$, $b\in\{0,1\}$, $Y\in\{\rightarrow,\leftarrow\}$,
$C^{b}_{i,Y}:\{0,1\}^{k}\times\{0,1\}^{r^{b}_{i}(n)}\rightarrow\{0,1\}^{k}\enspace.$
@@
which will exactly turn the state into $|{C_{1,\cdots,i+1}(R_{1,\cdots,i+1})}\rangle$. An additional way to look at this is that we want to apply a superposition over unitary choices, rather than orders, over the set $\{U_{j}\}_{j\in[m]}$, where we changed notation by $m=2^{\left(r_{i+1}\right)}$, $\forall j\in[m]:U_{j}:=U_{i+1,j}$. The amount of randomness $r_{i+1}\leq r(n)$ is logarithmic in $n$, thus the number of unitaries in the set $\{U_{j}\}_{j\in[m]}$ is polynomial in $n$ (and all of these unitaries are of polynomial time complexity). In order to finish our algorithm for state generation, it remains to only explain how to efficiently execute the above transformation of superposition of choice, with high probability.
@@
$\sum_{i\in[m]}U_{i}\cdot|{\psi}\rangle\enspace,$
with high probability. We wish to use OI transformation for this task. As in the case of OI transformations, we want the (pre-amplification) success probability of the transformation to be proportional to the norm of the generated state divided by the number of terms in the sum. In our case,
$\frac{\left\lVert\sum_{i\in[m]}U_{i}\cdot|{\psi}\rangle\right\rVert}{m}\enspace.$
Finally, before proceeding to solution, observe that being able to execute choice interference transformations with the above success probability solves the quantum state generation problem for $O(\log)$-sequentially invertible circuits, and thus overall $\mathsf{SISD}_{\mathsf{poly},O\left(\log(n)\right)}\in\textbf{BQP}^{\textbf{OI}}$.
@@
Generally when we execute OI transformation, the significant part of computational complexity is $m!$ divided by the norm of the OI state. As we explained earlier, the norm of the OI state ranges in $[0,m!]$, and the observation that leads our solution is the fact that there is a way to control the norm, by making unitary paths converge (while paying with limiting the variety of the superposition). For example, by taking all unitaries to be identical $U_{1}=\cdots=U_{m}$ the norm is $m!$, and the success probability is $1$. However, this is equivalent to simply executing $U_{1}^{m}$, which can be done by standard quantum computer.
Let $n\in\mathbb{N}$ the number of qubits in the state $|{\psi}\rangle$ (and the number of qubits that each of the unitaries $\{U_{i}\}_{i\in[m]}$ act on), and consider adding an ancilla of $\lceil\log{m}\rceil$ qubits, so the state we are operating on is $|{\psi}\rangle|{0^{\lceil\log{m}\rceil}}\rangle$. For each unitary $U_{i}$ we make a slight change, and execute the modified unitary $\tilde{U}_{i}$ which consists of two steps. (1) Execute the circuit $U^{c}_{i}$, which executes $U_{i}$ on the left $n$ qubits, conditioned on the state of the $\lceil\log{m}\rceil$ right qubits being $|{0^{\lceil\log{m}\rceil}}\rangle$. (2) Execute the unitary $U_{+1}$ on the right $\lceil\log{m}\rceil$ qubits, which treats the qubits as a binary representation of a number in $\{0,1,\cdots,m-1,\cdots,2^{\lceil\log{m}\rceil}-1\}$, and increments by $1$ modulo $2^{\lceil\log{m}\rceil}$. Here, we use the following fact, which we will also use later:
@@
For every numbers $a,N\in\mathbb{N}$ such that $a\leq N$, there is a unitary circuit $U_{a,N}$ on $\lceil\log(N)\rceil$ qubits, executing in time $\mathsf{poly}\left(\log\left(a\cdot N\right)\right)$ that maps
@@
That is, if we know the numbers $a,N$ not as part of the computation but in advance, then it is possible to unitarily and efficiently add $a$ modulo $N$. One can verify that the above fact follows from the function that adds $a$ modulo $N$ is both efficiently computable and invertible.
We consider what happens to the modified state $|{\psi}\rangle|{0^{\lceil\log{m}\rceil}}\rangle$, in each of the unitary execution orders $\sigma\in S_{m}$, for the modified unitaries $\tilde{U}_{i}$. Two things can be verified by the reader.
  • The first thing to see is that for $i\in[m]$ and any permutation $\sigma\in S_{m}$ such that $\sigma^{-1}\left(i\right)=1$, the state in the left $n$ qubits at the end of the computation is $U_{i}\cdot|{\psi}\rangle$. In simple terms, the only unitary that effectively executes on the side of the input state $|{\psi}\rangle$ is the first one in the ordering.
  • The second thing to notice is that when we look at the state of the right $\lceil\log(m)\rceil$-qubit counter register at the end of the computation, for any permutation $\sigma\in S_{m}$, the state is always $m\mod 2^{\lceil\log(m)\rceil}$.
Together, this means that the OI state is,
$\textbf{OI}\left(\{\tilde{U}_{i}\}_{i\in[m]},|{\psi}\rangle|{0^{\lceil\log(m)\rceil}}\rangle\right):=\sum_{\sigma\in S_{m}}\left(\prod_{i\in[m]}\tilde{U}_{\sigma^{-1}\left(i\right)}\right)\cdot|{\psi}\rangle|{0^{\lceil\log(m)\rceil}}\rangle$
$=\sum_{i\in[m]}\left(\sum_{\sigma\in S_{m-1}}\right)\left(U_{i}\cdot|{\psi}\rangle\right)\otimes|{m\text{ mod }2^{\lceil\log(m)\rceil}}\rangle$
$=\left(m-1\right)!\sum_{i\in[m]}\left(U_{i}\cdot|{\psi}\rangle\right)\otimes|{m\text{ mod }2^{\lceil\log(m)\rceil}}\rangle\enspace.$
So, if we get the normalized OI state, we can trace out the right $\lceil\log(m)\rceil$-qubit register without damaging the quantum information in the left $n$ qubits, and get the choice interference state. Finally, the scaled norm of the OI state is thus
$\frac{\left\lVert\textbf{OI}\left(\{\tilde{U}_{i}\}_{i\in[m]},|{\psi}\rangle|{0^{\lceil\log(m)\rceil}}\rangle\right)\right\rVert}{m!}$
$=\frac{\left\lVert\left(m-1\right)!\sum_{i\in[m]}\left(U_{i}\cdot|{\psi}\rangle\right)\otimes|{m\text{ mod }2^{\lceil\log(m)\rceil}}\rangle\right\rVert}{m!}$
$=\frac{\left\lVert\sum_{i\in[m]}\left(U_{i}\cdot|{\psi}\rangle\right)\otimes|{m\text{ mod }2^{\lceil\log(m)\rceil}}\rangle\right\rVert}{m}$
$=\frac{\left\lVert\sum_{i\in[m]}U_{i}\cdot|{\psi}\rangle\right\rVert}{m}\enspace,$
which concludes the proof.

### v2

@@ 2.1 Computable Order Interference and Quantum State Generation
When we think of the ability to generate new quantum states, there comes to mind one of the most fundamental hard problems in computer science – the Statistical Difference ($\mathsf{SD}$) problem [SV03]. As an intuitive description, the input to the $\mathsf{SD}$ problem is a pair of classically samplable distributions $D_{0}$, $D_{1}$, and the job of a solving algorithm is to distinguish whether the total variation distance between the distributions is small or large. Formally, for functions $a,b:\mathbb{N}\rightarrow[0,1]$ such that $\forall n\in\mathbb{N}:a(n)\leq b(n)$, an input to the $\mathsf{SD}_{a(n),b(n)}$ problem is a pair $\left(C_{0},C_{1}\right)$ of classical circuits with the same output size, that is, $C_{b}:\{0,1\}^{k_{b}}\rightarrow\{0,1\}^{k}$ for some $k_{0},k_{1},k\in\mathbb{N}$. For each of the two circuits, we consider the output distribution $D_{b}$ which is generated by executing $C_{b}$ on a uniformly random input $z\in\{0,1\}^{k_{b}}$. The problem comes with the promise that it is either the case that the total variation distance between $D_{0}$, $D_{1}$, is at most $a(n)$, in which case we denote $\left(C_{0},C_{1}\right)\in\prod_{\text{YES}}$, or the distance is more than $b(n)$, in which case we denote $\left(C_{0},C_{1}\right)\in\prod_{\text{NO}}$. Generally speaking, the larger the gap $b(n)-a(n)$, the easier the problem. In this work, our focus is on the family of problems $\mathsf{SD}_{a(n),b(n)}$, for functions $a(n),b(n)$ such that there exists a polynomial $\mathsf{poly}(n)$ with $b(n)^{2}-2a(n)+a(n)^{2}\geq\frac{1}{\mathsf{poly}(n)}$. From hereon we drop the subscript $a(n)$, $b(n)$ from $\mathsf{SD}_{a(n),b(n)}$ and write $\mathsf{SD}_{\mathsf{poly}}$ to implicitly refer to the family of problems $\mathsf{SD}_{a(n),b(n)}$ for such functions $a(n)$, $b(n)$. The $\mathsf{SD}_{\mathsf{poly}}$ problem is a notoriously hard problem in computer science (also for quantum computers), and it is at least as hard as some of the widely studied computational problems in cryptography and in combinatorics, for example: Quadratic Residuosity, Lattice Isomorphism, Discrete Logarithm, Decisional Diffie-Hellman, Graph Isomorphism, the Gap Closest Vector Problem and the Gap Shortest Vector Problem.
The thing that connects $\mathsf{SD}_{\mathsf{poly}}$ and the ability to generate non-trivial quantum states is a quantum procedure called the Swap Test [BBD+97, BCWDW01]. It is observed in [ATS03] that the $\mathsf{SD}_{\mathsf{poly}}$ problem could be efficiently solved if we had the ability to efficiently generate the output distribution states of classical circuits. If given input $\left(C_{0},C_{1}\right)$ we could efficiently generate the normalized $k$-qubit quantum states,
@@
then we can execute a Swap Test on the pair of states $|{C_{0}\left(R\right)}\rangle$, $|{C_{1}\left(R\right)}\rangle$. A Swap Test approximates (with some inverse polynomial error) the inner product between the states as vectors, which in turn relates to the total variation distance between the output distributions $D_{0}$, $D_{1}$ (by using the relations between fidelity and total variation distance). Finally, as long as the gap $b(n)^{2}-2a(n)+a(n)^{2}$ is at least inverse polynomial, the error in the approximation of the Swap Test can be overcome and we can decide whether the pair of circuits is in $\prod_{\text{YES}}$ or $\prod_{\text{NO}}$. So, the $\mathsf{SD}_{\mathsf{poly}}$ problem can be reduced to the quantum state generation of $|{C_{0}\left(R\right)}\rangle$, $|{C_{1}\left(R\right)}\rangle$. With accordance to the hardness of the $\mathsf{SD}_{\mathsf{poly}}$ problem, there is no known way for a quantum computer to generate in polynomial time the output distribution state for a general classical circuit $C$.
@@ 2.2 The Sequentially Invertible Statistical Difference Problem
It is a standard fact in quantum computing that for a classical circuit $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ there is an $\left(k^{\prime}+k\right)$-qubit unitary $U_{x,C}$, of roughly the same computational complexity as $C$, that maps: $\forall x\in\{0,1\}^{k^{\prime}},y\in\{0,1\}^{k}:U_{x,C}\cdot|{x,y}\rangle=|{x,y\oplus C(x)}\rangle$. It is in particular easy to quantumly generate the $\left(k^{\prime}+k\right)$-qubit state
@@
Sequentially invertible distributions.
@@
Our next idea is to parameterize invertibility. Our parameterization technique does not rely on any non-classical computation and may be of independent interest. Note that sampler circuits are equivalent to randomized circuits, which take the form $C:\{0,1\}^{k}\times\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$, where the randomness $z\in\{0,1\}^{k^{\prime}}$ is only considered as auxiliary input, and together with it there’s a main input $x\in\{0,1\}^{k}$. We can thus consider sampler and randomized circuits interchangeably.
@@
Our definition of parameterized invertibility of a circuit is as follows. Let $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ a circuit. For $r\in\{0,1,\cdots,k^{\prime}\}$, we say that $D$ is $r$-sequentially invertible if there is an $r$-invertible circuit sequence that samples from the output distribution of $C$ [^7] The full definition of $\left(r,t,\ell,\varepsilon\right)$-sequential invertibility, which takes into consideration more parameters and not only $r$, is given in Section 4.3. For the needs for this overview, we give the minimal definitions.. An $r$-invertible circuit sequence on $k$ bits is a sequence of pairs of circuits $\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ such that for every $i\in[\ell]$,
  • For every $Y\in\{\rightarrow,\leftarrow\}$, the circuit $C_{i,Y}$ uses $r_{i}\leq r$ random bits and maps from $k$ to $k$ bits, that is,
    $C_{i,Y}:\{0,1\}^{k}\times\{0,1\}^{r_{i}}\rightarrow\{0,1\}^{k}\enspace.$
    $\forall z\in\{0,1\}^{r_{i}},x\in\{0,1\}^{k}:x=C_{i,\leftarrow}\left(C_{i,\rightarrow}\left(x;z\right);z\right)\enspace.$
    Note that this also implies that for each hard-wired randomness, the circuits act as permutations on the set $\{0,1\}^{k}$.
@@
We say that the sequence samples from the output distribution of $C$ if we have the following equivalence of distributions,
$\Bigl{\{}C(z)\;|\;z\leftarrow\{0,1\}^{k^{\prime}}\Bigr{\}}\equiv$
@@
$\biggl{\{}C_{\ell,\rightarrow}\left(\cdots C_{2,\rightarrow}\left(C_{1,\rightarrow}\left(0^{k};z_{1}\right);z_{2}\right)\cdots;z_{\ell}\right)\;\Big{|}\;z_{1}\leftarrow\{0,1\}^{r_{1}},\cdots,z_{\ell}\leftarrow\{0,1\}^{r_{\ell}}\biggr{\}}\enspace.$
@@
The intuition is that the smaller the parameter $r$ is, ”the more the circuit is invertible”. We give a few examples. For any invertible circuit $C:\{0,1\}^{k}\rightarrow\{0,1\}^{k}$, $C$ is 00-sequentially invertible (for $\ell=1$, that is, by a single pair $\left(C_{1,\rightarrow},C_{1,\leftarrow}\right)$). For any (arbitrary) function $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$, $C$ is $k^{\prime}$-sequentially invertible (again for $\ell=1$). The function $C_{\oplus}:\{0,1\}^{2\cdot k}\rightarrow\{0,1\}^{k}$, that for two strings $a,b\in\{0,1\}^{k}$ outputs their bit-wise logical XOR: $C_{\oplus}(a,b):=a\oplus b$, is $1$-sequentially invertible (by a sequence of $\ell=k$ pairs), and $C_{\land}:\{0,1\}^{2\cdot k}\rightarrow\{0,1\}^{k}$, that for two strings $a,b\in\{0,1\}^{k}$ outputs their bit-wise logical AND: $C_{\land}(a,b):=a\land b$, is $2$-sequentially invertible (again, by a sequence of $k$ pairs). We leave it as an instructive exercise for the reader to verify the above claims.
@@
where for each $b\in\{0,1\}$, $\left(C^{b}_{i,\rightarrow},C^{b}_{i,\leftarrow}\right)_{i\in[\ell]}$ is an $r(n)$-invertible circuit sequence on $k$ bits, where $n$ is the input size,
$n:=\bigg{|}\left(\left(C^{0}_{i,\rightarrow},C^{0}_{i,\leftarrow}\right)_{i\in[\ell]},\left(C^{1}_{i,\rightarrow},C^{1}_{i,\leftarrow}\right)_{i\in[\ell]}\right)\bigg{|}\enspace.$
@@
which will exactly turn the state into $|{C_{1,\cdots,i+1}(R_{1,\cdots,i+1})}\rangle$. An additional way to look at this is that we want to apply a superposition over unitary choices, rather than orders, over the set $\{U_{j}\}_{j\in[m]}$, where we changed notation by $m=2^{\left(r_{i+1}\right)}$, $\forall j\in[m]:U_{j}:=U_{i+1,j}$. The amount of randomness $r_{i+1}\leq r(n)$ is logarithmic in $n$, thus the number of unitaries in the set $\{U_{j}\}_{j\in[m]}$ is polynomial in $n$ (and all of these unitaries are of polynomial time complexity).
@@
$\sum_{i\in[m]}U_{i}\cdot|{\psi}\rangle\enspace.$
For the solution to suffice, we need the success probability of choice interference to behave as the success probability of order interference, in both components of the probability: (1) the scaled norm, and (2) the phase alignment. The scaled norm should be the state’s norm divided by the number of choices $m$, and the phases should align between choices, rather than all orders. Formally, we want to construct an efficient quantum algorithm that has access to the order interference oracle, and can execute an interference of choices with success probability of:
$\left(\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{i\in[m]}\alpha_{x,i}|}{\sum_{x\in\{0,1\}^{n}}\sum_{i\in[m]}|\alpha_{x,i}|}\right)\frac{\frac{\left\lVert\sum_{i\in[m]}U_{i}\cdot|{\psi}\rangle\right\rVert}{m}}{\frac{\left\lVert\sum_{i\in[m]}U_{i}\cdot|{\psi}\rangle\right\rVert}{m}+\frac{1}{\lambda}}\enspace,$ (8)
where for $i\in[m]$ we write,
$U_{i}\cdot|{\psi}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{x,i}\cdot|{x}\rangle\enspace,$
Lastly, for the specific type of unitaries that we use for the algorithm for state generation (i.e., unitaries that implement the action of invertible classical circuits), all amplitudes are real and non-negative in each step, and thus in particular phases perfectly align. One can observe that in order to finish our algorithm for state generation (which in turn will finish the proof for $\mathsf{SISD}_{\mathsf{poly},O\left(\log(n)\right)}\in\textbf{BQP}^{\textbf{OI}}$), it remains to only explain how to efficiently execute the above transformation with the target success probability.
@@
The observation at the center of our solution is that we can make unitary paths converge, while paying with limiting the variety of the superposition. For example, by taking all unitaries to be identical $U_{1}=\cdots=U_{m}$ the norm of the OI state is $m!$ and the phases of all unitary evolution orders align, which makes the success probability $1$. However, this is equivalent to simply executing $U_{1}^{m}$, which can be done by a standard quantum computer. Our solution to this is as follows.
Let $n\in\mathbb{N}$ the number of qubits in the state $|{\psi}\rangle$ (and the number of qubits that each of the unitaries $\{U_{i}\}_{i\in[m]}$ act on), and consider adding an ancilla of $\lceil\log{m}\rceil$ qubits, so the state we are operating on is $|{\psi}\rangle|{0^{\lceil\log{m}\rceil}}\rangle$. For each unitary $U_{i}$ we make a slight change, and execute the modified unitary $\tilde{U}_{i}$ which consists of two steps. (1) Execute the circuit $U^{c}_{i}$, which executes $U_{i}$ on the left $n$ qubits, conditioned on the state of the $\lceil\log{m}\rceil$ right qubits being $|{0^{\lceil\log{m}\rceil}}\rangle$. (2) Execute the unitary $U_{+1}$ on the right $\lceil\log{m}\rceil$ qubits, which treats the qubits as a binary representation of a number in $\{0,1,\cdots,m-1,\cdots,2^{\lceil\log{m}\rceil}-1\}$, and increments by $1$ modulo $2^{\lceil\log{m}\rceil}$. For the Unitary $U_{+1}$ we use the following fact, which we will also use later:
@@
For any pair of numbers $a,N\in\mathbb{N}$ such that $a\leq N$, the function $f_{a,N}:\{0,1\}^{\lceil\log(N)\rceil}\rightarrow\{0,1\}^{\lceil\log(N)\rceil}$ that adds to the input $a$ modulo $N$ is efficiently computable and invertible. It follows there is a unitary circuit $U_{a,N}$ on $\lceil\log(N)\rceil$ qubits, executing in time $\mathsf{poly}\left(\log\left(a\cdot N\right)\right)$ that maps
@@
That is, if we know the numbers $a,N$ not as part of the computation but in advance, then it is possible to efficiently add $a$ modulo $N$, in superposition. The solution to executing choice interference is to execute a call to the OI oracle with input ($\left(\{\tilde{U}_{i}\}_{i\in[m]},\;|{\psi}\rangle|{0^{\lceil\log{m}\rceil}}\rangle\right)$, and then tracing out the rightmost $\lceil\log{m}\rceil$ qubits.
Let us analyze the result of the above suggestion. We consider what happens to the modified state $|{\psi}\rangle|{0^{\lceil\log{m}\rceil}}\rangle$, in each of the unitary execution orders $\sigma\in S_{m}$, for the modified unitaries $\tilde{U}_{i}$. Two things can be verified by the reader.
  • For $i\in[m]$ and any permutation $\sigma\in S_{m}$ such that $\sigma^{-1}\left(i\right)=1$, the state in the left $n$ qubits at the end of the computation is $U_{i}\cdot|{\psi}\rangle$. In simple terms, the only unitary that effectively executes on the side of the input state $|{\psi}\rangle$ is the first one in the ordering $\sigma$.
  • When we look at the state of the right $\lceil\log(m)\rceil$-qubit counter register at the end of the computation, for any execution order $\sigma\in S_{m}$, the state is always $m\mod 2^{\lceil\log(m)\rceil}$. This in particular means that the rightmost $\lceil\log{m}\rceil$ qubits are disentangled from the rest of the left $n$ qubits, and can be traced out without damaging coherence of the rest of the state.
Together, this also means that for every $i\in[m]$,
$\sum_{\sigma\in S_{m}:\sigma^{-1}(1)=i}\left(\prod_{j\in[m]}U_{\sigma^{-1}\left(j\right)}\right)\cdot|{\psi}\rangle=\left(m-1\right)!\cdot\left(U_{i}\cdot|{\psi}\rangle\right)\otimes|{m\text{ mod }2^{\lceil\log(m)\rceil}}\rangle\enspace.$
Finally, since all $m!$ permutations in $S_{m}$ can be partitioned into $m$ sets of equal size $\left(m-1\right)!$, such that for $i\in[m]$, set $i$ is all permutations $\sigma\in S_{m}$ such that $\sigma^{-1}(1)=i$, by calculation, we get exactly the desired success probability in 8.